OpenStudy (minime):

2x+5y= 2 3x-2y= -16 please solve using elmination

OpenStudy (minime):

helpp

OpenStudy (lalaly):

multiply first by 3 and sec by 2 6x+15y=6 6x-4y=-32 -------------------------now subtract 0+19y=38 y=2

OpenStudy (lalaly):

then substitute y=2 in any equation to find x

OpenStudy (minime):

where did the 15 come from

OpenStudy (lalaly):

i multiplied the first equation by3 to make the x coeffecient of both equations the same...so u can cancel it out in subtraction

OpenStudy (minime):

shutup

OpenStudy (lalaly):

:S